# JanusFlow: Harmonizing Autoregression and Rectified Flow for Unified Multimodal Understanding and Generation

### v1

@@ 1 Introduction
JanusFlow shows state-of-the-art performances in both multimodal comprehension and text-to-image generation compared to existing unified approaches, and even outperforms several specialized methods. Specifically, on text-to-image generation benchmarks, MJHQ FID-3030k [48], GenEval [28] and DPG-Bench [34], JanusFlow achieves scores of $9.51$, $0.63$ and $80.09\%$, surpassing established text-to-image models including SDv1.5 [75] and SDXL [71]. In multimodal comprehension benchmarks, JanusFlow attains scores of $74.9$, $70.5$ and $60.3$ on MMBench [62], SeedBench [46], and GQA [35], respectively, exceeding specialized models such as LLaVA-v1.5 [56] and Qwen-VL-Chat [4]. Notably, these results are achieved with a compact LLM architecture with only 1.3B parameters.
@@ 2 Related Work
Our work focuses on developing unified models that combine autoregressive capabilities with flow/diffusion models, leveraging their proven effectiveness in visual generation. Compared to similar approaches [96, 103], JanusFlow offers three key advantages: (i) a simple yet effective generation process using rectified flow, (ii) enhanced performance through decoupled vision encoders that resolve inter-task conflicts, and (iii) improved generation quality through representation alignment regularization, enabled by our decoupled encoder design.
@@ 3.1 Background
$\frac{{\rm{d}}z_{t}}{{\rm{d}}t}=v_{\theta_{NN}}(z_{t},t),\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ z_{0}\sim\pi_{0},$ (2)
@@
[…]
@@ 3.2 A Unified Framework for Multimodal Understanding and Generation
$v(z_{t},t)=wv(z_{t},t\penalty 10000\ |\penalty 10000\ x^{con})+(1-w)v(z_{t},t\penalty 10000\ |\penalty 10000\ \varnothing),$ (5)
where $v(z_{t},t\penalty 10000\ |\penalty 10000\ \varnothing)$ denotes the velocity inferred without text conditioning and $w\geqslant 1$ controls the magnitute of CFG. Empirically, increasing $w$ yields higher semantic alignment [75, 61, 71, 23]. Analogous to multimodal understanding, we prepend the special token |BOI| to indicate the start of image generation in the sequence.
@@ 3.4 Training Objective
Training JanusFlow involves two types of data, multimodal understanding data and image generation data. Both types of data contain two parts: “condition” and “response”. “Condition” refers to the prompting of the tasks (e.g., text prompts in the task of generation and images in the task of understanding) while “response” refers to the corresponding responses of the two tasks. The data can be formatted as $x=(x^{con},x^{res})$, where the superscript concon denotes “condition” and resres denotes “response”. We denote the length of the whole sequence $x$ as $\ell$, the length of $x^{con}$ as $\ell_{con}$ and the length of $x^{res}$ as $\ell_{res}$. We use $\theta$ to represent the collection of all the trainable parameters in JanusFlow, including the LLM, $f_{enc}$, $g_{enc}$, $g_{dec}$ and the linear transformation layers.
@@
$\mathcal{L}_{RF}(\theta)={{\mathbb{E}}}_{x\sim\mathcal{D}_{gen},t\sim{{\rm{P}}}(t),z_{0}\sim\mathcal{N}(0,I)}\left[\left|\left|v_{\theta}(z_{t},t\penalty 10000\ |\penalty 10000\ x^{con})-(x^{res}-z_{0})\right|\right|^{2}\right],$ (7)
@@
Learning Rate $0\times 10^{-4} 1×1$$\times 10^{-4}$ $2.0\times 10^{-5}$
@@
Warm-up Steps $2,000$ 0 $1,000$
Training Steps $10,000$ $380,000$ $26,000$
@@ 4.2 Training Data Settings
    Multimodal Understanding Data. This type of data contains several sub-categories: (a) Image caption data. We incorporate caption datasets from [20, 41, 50, 51, 53, 79] and generate additional captions for images from [16, 43] using open-source multimodal understanding models. The data follows template formats, e.g., “<image>Generate the caption of this picture. <caption>”. (b) Charts and tables. We directly adopt the chart and table data from the training data of DeepSeek-VL [63]. (c) Task data. ShareGPT4V [11] data is utilized to facilitate basic question-answering capabilities during pre-training, structured as “<image><question><answer>”. (d) Interleaved text-image data. This sub-category is sourced from [81, 42].
@@
Gen. Only LlamaGen [83] 0.80.8B $0.71$ $0.34$ $0.21$ $0.58$ $0.07$ $0.04$ $0.32$
LDM [75] 1.41.4B $0.92$ $0.29$ $0.23$ $0.70$ $0.02$ $0.05$ $0.37$
SDv1.51.5 [75] 0.90.9B $0.97$ $0.38$ $0.35$ $0.76$ $0.04$ $0.06$ $0.43$
PixArt-$\alpha$ [9] 0.60.6B $0.98$ $0.50$ $0.44$ $0.80$ $0.08$ $0.07$ $0.48$
SDv2.12.1 [75] 0.90.9B $0.98$ $0.51$ $0.44$ $0.85$ $0.07$ $0.17$ $0.50$
DALL-E $2$ [74] 6.56.5B $0.94$ $0.66$ $0.49$ $0.77$ $0.10$ $0.19$ $0.52$
Emu33-Gen [91] 88B $0.98$ $0.71$ $0.34$ $0.81$ $0.17$ $0.21$ $0.54$
SDXL [71] 2.62.6B $0.98$ $0.74$ $0.39$ $0.85$ $0.15$ $0.23$ $0.55$
IF-XL [17] 4.34.3B $0.97$ $0.74$ $0.66$ $0.81$ $0.13$ $0.35$ $0.61$
@@
Unified Chameleon [85] 3434B - - - - - - $0.39$
LWM [58] 77B $0.93$ $0.41$ $0.46$ $0.79$ $0.09$ $0.15$ $0.47$
SEED-X [27] 1717B $0.97$ $0.58$ $0.26$ $0.80$ $0.19$ $0.14$ $0.49$
Show-o [96] 1.31.3B $0.95$ $0.52$ $0.49$ $0.82$ $0.11$ $0.28$ $0.53$
Janus [93] 1.31.3B $0.97$ $0.68$ $0.30$ $0.84$ $0.46$ $0.42$ $0.61$
@@ 4.3 Evaluation Settings
We evaluate JanusFlow’s multimodal understanding abilities across a diverse set of vision-language benchmarks for general understanding capabilities, including POPE [52], MME [24], MMBench [62], SEEDBench [46], VQAv2 [29], GQA [35], $M$-Vet [100], and MMMU [101].
@@ 4.4 Quantitative Results
Und. Only MobileVLM [12] 2.72.7B $84.9$ $1288.9$ $59.6$ - - $59.0$ - -
MobileVLM-V2 [13] 2.72.7B $84.7$ $1440.5$ $63.2$ - - $61.1$ - -
LLaVA-Phi [104] 2.72.7B $85.0$ $1335.1$ $59.8$ - $71.4$ - - $28.9$
LLaVA [57] 77B $76.3$ $809.6$ $38.7$ $33.5$ - - - $25.5$
LLaVA-v1.51.5 [56] 77B $85.9$ $1510.7$ $64.3$ $58.6$ $78.5$ $62.0$ $35.4$ $31.1$
InstructBLIP [15] 77B - - $36.0$ $53.4$ - $49.2$ - $26.2$
Qwen-VL-Chat [4] 77B - $1487.5$ $60.6$ $58.2$ $78.2$ $57.5$ - -
IDEFICS-99B [44] 88B - - $48.2$ - $50.9$ $38.4$ - -
Emu33-Chat [91] 88B $85.2$ - $58.5$ $68.2$ $75.1$ $60.3$ $31.6$ -
InstructBLIP [15] 1313B $78.9$ $1212.8$ - - - $49.5$ - $25.6$
LLaVA-v1.51.5-Phi-$1.5$ [96] 1.31.3B $84.1$ $1128.0$ - - $75.3$ $56.5$ $30.7$ -
MobileVLM [12] 1.41.4B $84.5$ $1196.2$ $53.2$ - - $56.1$ - -
MobileVLM-V2 [13] 1.41.4B $84.3$ $1302.8$ $57.7$ - - $59.3$ - -
Unified Gemini-Nano-1 [86] 1.81.8B - - - - $62.7$ - $26.3$ -
LWM [58] 77B $75.2$ - - - $55.8$ $44.8$ - $9.6$
VILA-U [95] 77B $85.8$ $1401.8$ - $59.0$ $79.4$ $60.8$ - $33.5$
Chameleon [85] 77B - - - - - - $22.4$ $8.3$
DreamLLM [19] 77B - - - - $72.9$ - - $36.6$
LaVIT [37] 77B - - - - $66.0$ $46.8$ - -
Emu [84] 1313B - - - - $52.0$ - - -
NExT-GPT [94] 1313B - - - - $66.7$ - - -
Show-o [96] 1.31.3B $73.8$ $948.4$ - - $59.3$ $48.7$ $25.1$ -
Janus [93] 1.31.3B $87.0$ $1338.0$ $69.4$ $63.7$ $77.3$ $59.1$ $30.5$ $34.3$
JanusFlow (Ours) 1.3B 88.0 1333.1 74.9 70.5 79.8 60.3 29.3 30.9
@@ Appendix A Performance Analysis of 256 Resolution Model
Appendix B Analysis of CFG Factor and Sampling Steps
@@
Appendix C Additional Qualitative Results
Figure 2: More multimodal understanding cases.
Figure 3: More text-to-image generation results.

### v2

@@ 1 Introduction
JanusFlow shows state-of-the-art performances in both multimodal comprehension and text-to-image generation compared to existing unified approaches, and even outperforms several specialized methods. Specifically, on text-to-image generation benchmarks, MJHQ FID-$30$k [48], GenEval [28] and DPG-Bench [34], JanusFlow achieves scores of $9.51$, $0.63$ and $80.09\%$, surpassing established text-to-image models including SDv1.5 [77] and SDXL [73]. In multimodal comprehension benchmarks, JanusFlow attains scores of $74.9$, $70.5$ and $60.3$ on MMBench [63], SeedBench [46], and GQA [35], respectively, exceeding specialized models such as LLaVA-v1.5 [56] and Qwen-VL-Chat [4]. Notably, these results are achieved with a compact LLM architecture with only 1.3B parameters.
@@ 2 Related Work
Our work focuses on developing unified models that combine autoregressive capabilities with flow/diffusion models, leveraging their proven effectiveness in visual generation. Compared to similar approaches [100, 108, 107], JanusFlow offers three key advantages: (i) a simple yet effective generation process using rectified flow, (ii) enhanced performance through decoupled vision encoders that resolve inter-task conflicts, and (iii) improved generation quality through representation alignment regularization, enabled by our decoupled encoder design.
@@ 3.1 Background
$\frac{{\rm{d}}z_{t}}{{\rm{d}}t}=v_{\theta_{NN}}(z_{t},t),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ z_{0}\sim\pi_{0},$ (2)
@@
$\min_{\theta}{{\mathbb{E}}}_{t\sim{{\rm{P}}}(t),z_{0}\sim\pi_{0},x\sim\pi_{1}}\left[\left|\left|v_{\theta_{NN}}(z_{t},t)-(x-z_{0})\right|\right|^{2}\right],\leavevmode\nobreak\ \leavevmode\nobreak\ \text{where}\leavevmode\nobreak\ \leavevmode\nobreak\ z_{t}=tx+(1-t)z_{0}.$ (3)
@@ 3.2 A Unified Framework for Multimodal Understanding and Generation
$v(z_{t},t)=wv(z_{t},t\leavevmode\nobreak\ |\leavevmode\nobreak\ x^{con})+(1-w)v(z_{t},t\leavevmode\nobreak\ |\leavevmode\nobreak\ \varnothing),$ (5)
where $v(z_{t},t\leavevmode\nobreak\ |\leavevmode\nobreak\ \varnothing)$ denotes the velocity inferred without text conditioning and $w\geqslant 1$ controls the magnitute of CFG. Empirically, increasing $w$ yields higher semantic alignment [77, 62, 73, 23]. Analogous to multimodal understanding, we prepend the special token |BOI| to indicate the start of image generation in the sequence.
@@ 3.4 Training Objective
Training JanusFlow involves two types of data, multimodal understanding data and image generation data. Both types of data contain two parts: “condition” and “response”. “Condition” refers to the prompting of the tasks (e.g., text prompts in the task of generation and images in the task of understanding) while “response” refers to the corresponding responses of the two tasks. The data can be formatted as $x=(x^{con},x^{res})$, where the superscript $con$ denotes “condition” and $res$ denotes “response”. We denote the length of the whole sequence $x$ as $\ell$, the length of $x^{con}$ as $\ell_{con}$ and the length of $x^{res}$ as $\ell_{res}$. We use $\theta$ to represent the collection of all the trainable parameters in JanusFlow, including the LLM, $f_{enc}$, $g_{enc}$, $g_{dec}$ and the linear transformation layers.
@@
$\mathcal{L}_{RF}(\theta)={{\mathbb{E}}}_{x\sim\mathcal{D}_{gen},t\sim{{\rm{P}}}(t),z_{0}\sim\mathcal{N}(0,I)}\left[\left|\left|v_{\theta}(z_{t},t\leavevmode\nobreak\ |\leavevmode\nobreak\ x^{con})-(x^{res}-z_{0})\right|\right|^{2}\right],$ (7)
@@
Learning Rate $1.0\times 10^{-4}$ $1\times 10^{-4}$ $2.0\times 10^{-5}$
@@
Warm-up Steps $2,000$ $2,000$ $1,000$
Training Steps $10,000$ $390,000$ $26,000$
@@ 4.2 Training Data Settings
    Multimodal Understanding Data. This type of data contains several sub-categories: (a) Image caption data. We incorporate caption datasets from [20, 41, 50, 51, 53, 82] and generate additional captions for images from [16, 43] using open-source multimodal understanding models. The names of the datasets are provided in the supplementary materials. The data follows template formats, e.g., “<image>Generate the caption of this picture. <caption>”. (b) Charts and tables. We directly adopt the chart and table data from the training data of DeepSeek-VL [64]. (c) Task data. ShareGPT4V [11] data is utilized to facilitate basic question-answering capabilities during pre-training, structured as “<image><question><answer>”. (d) Interleaved text-image data. This sub-category is sourced from [84, 42].
@@
Gen. Only LlamaGen [86] $0.8$B $0.71$ $0.34$ $0.21$ $0.58$ $0.07$ $0.04$ $0.32$
LDM [77] $1.4$B $0.92$ $0.29$ $0.23$ $0.70$ $0.02$ $0.05$ $0.37$
SDv1.51.51.51.5 [77] $0.9$B $0.97$ $0.38$ $0.35$ $0.76$ $0.04$ $0.06$ $0.43$
PixArt-$\alpha$ [9] $0.6$B $0.98$ $0.50$ $0.44$ $0.80$ $0.08$ $0.07$ $0.48$
SDv2.12.12.12.1 [77] $0.9$B $0.98$ $0.51$ $0.44$ $0.85$ $0.07$ $0.17$ $0.50$
DALL-E $2$ [76] $6.5$B $0.94$ $0.66$ $0.49$ $0.77$ $0.10$ $0.19$ $0.52$
Emu3333-Gen  [95] $8$B $0.98$ $0.71$ $0.34$ $0.81$ $0.17$ $0.21$ $0.54$
SDXL [73] $2.6$B $0.98$ $0.74$ $0.39$ $0.85$ $0.15$ $0.23$ $0.55$
IF-XL [17] $4.3$B $0.97$ $0.74$ $0.66$ $0.81$ $0.13$ $0.35$ $0.61$
@@
Unified Chameleon [88] $34$B - - - - - - $0.39$
LWM [59] $7$B $0.93$ $0.41$ $0.46$ $0.79$ $0.09$ $0.15$ $0.47$
SEED-X [27] $17$B $0.97$ $0.58$ $0.26$ $0.80$ $0.19$ $0.14$ $0.49$
Show-o [100] $1.3$B $0.95$ $0.52$ $0.49$ $0.82$ $0.11$ $0.28$ $0.53$
Janus [97] $1.3$B $0.97$ $0.68$ $0.30$ $0.84$ $0.46$ $0.42$ $0.61$
Transfusion [108] $7.3$B - - - - - - 0.63
@@ 4.3 Evaluation Settings
We evaluate JanusFlow’s multimodal understanding abilities across a diverse set of vision-language benchmarks for general understanding capabilities, including POPE [52], MME [24], MMBench [63], SEEDBench [46], VQAv2 [29], GQA [35], MM-Vet [104], MMMU [105], ChartQA[70] and TextVQA[81]
@@ 4.4 Quantitative Results
Type Model LLM Param POPE MME-P MMBdev SEED VQAv2test GQA MMMU MM-Vet ChartQA TextVQA
Und. Only MobileVLM [12] $2.7$B $84.9$ $1288.9$ $59.6$ $59.0$ $47.5$
MobileVLM-V2 [13] $2.7$B $84.7$ $1440.5$ $63.2$ $61.1$ $57.5$
LLaVA-Phi [109] $2.7$B $85.0$ $1335.1$ $59.8$ $71.4$ $28.9$ $48.6$
LLaVA [58] $7$B $76.3$ $809.6$ $38.7$ $33.5$ $25.5$
LLaVA-v1.5 [56] $7$B $85.9$ $1510.7$ $64.3$ $58.6$ $78.5$ $62.0$ $35.4$ $31.1$ $58.2$
InstructBLIP [15] $7$B - - $36.0$ $53.4$ $49.2$ $26.2$ $50.1$
Qwen-VL-Chat [4] $7$B - $1487.5$ $60.6$ $58.2$ $78.2$ $57.5$ $66.3$ $61.5$
LLaVA-NeXT [57] $7$B - $1519.3$ $35.1$ $54.8$
Qwen2-VL [94] $7$B - - - - - - $54.1$ $62.0$ $83.0$ $84.3$
IDEFICS-$9$[44] $8$B - - $48.2$ $50.9$ $38.4$ $25.9$
Emu3333-Chat [95] $8$B $85.2$ $58.5$ $68.2$ $75.1$ $60.3$ $31.6$ $68.6$ $64.7$
InstructBLIP [15] $13$B $78.9$ $1212.8$ $49.5$ $25.6$ $50.7$
@@
LLaVA-v1.5-Phi-1.5 [100] $1.3$B $84.1$ $1128.0$ $75.3$ $56.5$ $30.7$
MobileVLM [12] $1.4$B $84.5$ $1196.2$ $53.2$ $56.1$ $41.5$
MobileVLM-V2 [13] $1.4$B $84.3$ $1302.8$ $57.7$ $59.3$ $52.1$
Unified Gemini-Nano-1 [89] $1.8$B - - - - $62.7$ $26.3$ $53.6$ $62.5$
LWM [59] $7$B $75.2$ $55.8$ $44.8$ $9.6$
VILA-U [99] $7$B $85.8$ $1401.8$ $59.0$ $79.4$ $60.8$ $33.5$ $60.8$
Chameleon [88] $7$B - - - - - - $22.4$ $8.3$
DreamLLM [19] $7$B - - - - $72.9$ $36.6$ $41.8$
LaVIT [37] $7$B - - - - $66.0$ $46.8$
Emu [87] $13$B - - - - $52.0$
NExT-GPT [98] $13$B - - - - $66.7$
[…]
Show-o [100] $1.3$B $73.8$ $948.4$ $59.3$ $48.7$ $25.1$
Janus [97] $1.3$B $87.0$ $1338.0$ $69.4$ $63.7$ $77.3$ $59.1$ $30.5$ $34.3$
JanusFlow (Ours) $1.3$B $88.0$ $1333.1$ $74.9$ $70.5$ $79.8$ $60.3$ $29.3$ $30.9$ $64.6$ $55.5$
@@ Appendix A Performance Analysis of 256 Resolution Model
Appendix B Details of the Datasets
The datasets used in the pre-training stage for understanding include DetailedCaption [20], SAM [41], arXivQA[50], DenseFusion-1M [51], MMSci[53], PixelProse [82], re-captioned LAION-Aesthetics [16], re-captioned Open Images V4 [43], ShareGPT4V [11], WikiHow [42] and WIT [84]. The datasets used in the pre-training stage for generation include re-captioned LAION-Aesthetics [16], DALL-E 3 1M [21], SAM [41], Open Images V4 [43], Megalith-10M [68], YFCC-15M [71], PixelProse[82] and JourneyDB [85].
Appendix C Analysis of CFG Factor and Sampling Steps
@@ Appendix D Details of REPA Ablation
Appendix E Additional Qualitative Results
Figure 3: More multimodal understanding cases.
Figure 4: More text-to-image generation results.